# Segment Anything Model for Road Network Graph Extraction

### v1

@@ 1 Introduction
Road network graphs are spatial representations of the structure and layout of road networks. They are typically stored in a vectorized format, consisting of vertices and edges. The vertices may represent intersections, and edges could stand for road segments. Large-scale road network graphs are vital for various applications: they enable navigation systems like Google Maps to determine optimal routes, assist in path planning for autonomous vehicles, and help city planners in traffic analysis and optimization, to name a few. These applications call for accurate and efficient methods to automatically create such graphs, as they require scaling to huge regions and even near-continuous updating, which are astoundingly expensive when manually done. Therefore, systems for automatically generating such maps have tremendous application value and are under active research.
@@
Recently, the rapid growth of foundational models has showcased their impressive capabilities. These models, which leverage flexible, high-capacity, and scalable architectures such as Transformers, are pre-trained through effective self-supervision methods and unprecedentedly large datasets. This pre-training endows them with robust semantic reasoning and generalizability. An example of such models is the Segment Anything Model (SAM) [16], which, built on the powerful ViT [8] architecture and trained with millions to billions of images and masks, demonstrates unparalleled semantic segmentation capabilities. This raises intriguing questions: How can SAM be applied to the prediction of road network graphs from satellite images, and how good can it be?
@@
In this work, we answer these questions by introducing the SAM-Road model, which adapts the SAM [16] for generating large-scale, vectorized road network graphs. This adaptation incorporates domain knowledge from previous research in satellite mapping. We divide the problem into two main components: geometric prediction and topology reasoning.
@@
We model graph geometry with a set of 2D vertices that, when densely sampled, accurately reflect the graph’s overall shape. Drawing on foundational research in satellite mapping, we approach geometric prediction as a dense semantic segmentation task. The SAM-Road model predicts dense masks to indicate the likelihood of road elements such as lane segments and intersections, then employs simple non-maximum suppression to convert the pixels into vertices of the desired density. Leveraging the inherent semantic segmentation capabilities of SAM [16], this method can effectively capture highly complex shapes, which are common in dense urban areas.
@@
A notable challenge for segmentation-based mapping approaches is the difficulty of inferring topology from dense imagery. This branch of methods often relies on slow and complex post-processing heuristics, and they are still error-prone even with additional channels indicating orientation and connectivity. Inspired by recent advances in learning-based graph prediction, we developed a transformer-based graph neural network as the second stage of our model, following SAM segmentation. This network focuses on predicting the local subgraph around each vertex and determining connectivity with nearby vertices to establish the overall graph topology. It utilizes relative vertex positions and image embeddings from the SAM backbone to guide its predictions.
@@
Despite its straightforward design, SAM-Road achieves accuracy comparable to more complex state-of-the-art systems on two widely recognized satellite mapping datasets: City-scale [12] and SpaceNet [29]. Moreover, for large spatial areas spanning multi-square kilometers, its architecture supports high degrees of parallelism and rapid GPU inference, achieving speeds up to 80 times faster than existing methods. We hope that this work will inspire further exploration of foundational vision models in remote sensing and graph learning tasks. Code will be made available to support future research.
@@ 2 Related works
2.1 SAM and applications
In 2023, the Segment Anything paper [16] proposed a foundational model for image segmentation, capable of generalizing across various tasks, enabling impressive zero-shot transfer capabilities. Through fine-tuning or direct adoption, SAM has been used in object detection [20, 7], image impainting [39], segmentation of medical images [23, 15, 32], and remote sensing tasks [5]. Existing adaptations of SAM [16] in remote sensing have focused more on simple segmentation and have not been applied to the production of road network graphs at the time of this work.
@@ 2.2 Road network graph prediction
Segmentation-based methods [36, 37, 2, 1, 12, 24] treat the task as a dense mask prediction. They represent the road network graph structure through one or more images, each detailing aspects such as road existence, intersections [36], orientation [2], and connectivity [12]. Post-processing heuristics, such as thinning [10] and path-finding [1], are then employed to extract the vectorized graph structure. Benefits of this approach include 1) the ability of segmentation masks to represent complex geometries as a bottom-up volumetric representation [13], and 2) ease of parallel patch-wise inference for large areas, and subsequent result aggregation for refinement. However, the challenge of topology prediction persists: handcrafted heuristics often fail with poor mask quality; even with high-quality masks, deriving topology from them remains ill-formed. There exist no universal heuristics for all complex road structures, like multi-road intersections, multi-lane highways, and overpasses. Moreover, the heuristic tends to rely on CPU-intensive logic, which often becomes the inference speed bottleneck.
@@ 2.3 Graph representation and learning
Graph representation and learning involves mapping data to graph structures and applying learning algorithms to understand complex relationships within the data. Graphs are composed of nodes (vertices) and edges (links between nodes) [11, 30, 35], making them ideal for modeling complex systems with interrelated elements, such as social networks, molecular structures, and communication networks [31, 28].
Significant advancements have been seen in this area with the development of Graph Neural Networks (GNNs) [34], Graph Convolutional Networks (GCNs) [17], and Transformers [8] adapted for graph data. GNNs offer powerful mechanisms for aggregating neighborhood information and understanding node relationships, making them suitable for road network analysis. GCNs, by applying convolution over graphs to capture local node features within road networks, can aid in graph extraction tasks. Transformers, through self-attention mechanisms, have been adapted to capture long-range dependencies within graphs, enhancing connectivity prediction and scene graph generation.
Our method combines the advantages of segmentation-based and graph-based approaches. It harnesses the exceptional capabilities of SAM [16] to generate a high-quality mask for geometry prediction, and uses a transformer-based graph neural network to directly produce graph structures without handcrafted post-processing heuristics.
@@ 3.1 Overall Architecture
Figure 1: The architecture of our approach, SAM-Road. It contains an image encoder taken from the pre-trained SAM [16], a geometry decoder, and a topology decoder. It directly predicts vectorized graph vertices (yellow) and edges (orange) from an input RGB satellite imagery.
The overall structure of SAM-Road is shown in Figure  1. It contains an image encoder taken from the pre-trained SAM [16], a geometry decoder, and a topology decoder. The model takes as input an RGB satellite imagery. First, the image encoder produces the image feature embeddings. Then, the geometry decoder predicts the per-pixel existence probability, for both roads and intersections. The set of graph vertices $\mathbf{V}\{v_{i}\in\mathbb{R}^{2}\}$ representing 2D locations is extracted from these masks with a simple non-maximum suppression process, detailed in Algorithm 1. Given the predicted vertices, the topology decoder goes over each of them and determines whether it should connect to its nearby vertices within a given radius $R_{\text{nbr}}$ , given its local context. For an edge $(v_{i},v_{j})$, it predicts the probability that it exists. One edge may be predicted more than once, its final score will be the average. Eventually, the road network graph $\mathbf{G}$ is predicted as the sets of vertex $\mathbf{V}$ and edges $\mathbf{E}$
@@ 3.3 Geometry Decoder
We predict masks for both intersections and roads for more accurate graph structures at intersections. If only the road mask existed, there would be no guarantee that the center point of an intersection would be kept, producing error patterns like Figure  5. To mitigate this: 1) Vertices are extracted from both masks with the same NMS algorithm. 2) The two sets of vertices are joined, with all intersection vertices assigned a higher score than any road vertices. 3) The joined set is then NMS-ed again to produce the final result. This ensures intersection points are kept as much as possible.
@@ 3.4 Topology Decoder
The connection here is defined as ”whether two vertices are immediate neighbors on the graph”. That is, imagine a breadth-first-search on the road network graph from the source vertex, which stops expanding whenever a) it hits a target vertex or b) the depth (search radius) exceeds $R_{\text{nbr}}$ a target vertex is only connected to the source if it’s visited by the search. This is further illustrated in Figure  2.
@@ 3.5 Label generation
Mask Labels. For road mask labels, we rasterize the ground-truth road lines, by drawing each edge as a line segment, with a width of 3 pixels. The pixels covered by the line segments are set to 1, and others are 0. For intersection labels, we find all the graph vertices with a degree not equal to 2 and render them as circles with a radius of 3 pixels. This is partially inspired by the OpenPose work which represents human keypoint graphs as heatmaps.
Topology Labels. During training, we don’t run the vertex extraction process. The topology decoder is trained in a teacher-forcing manner, where the vertices being asked are not from model prediction, but sampled from ground-truth road network graphs to emulate the predictions. This is done by first subdividing the ground-truth graph and then running the same NMS procedure as the inference stage. To emulate various NMS results, a uniform random score is assigned to each subdivision vertex.
@@ 3.6 Sliding-window Inference for Large Regions
Figure 3: SAM-Road can predict the entire road network graph for arbitrarily large regions by operating in a sliding-window manner. It first extracts the global nodes, caches the per-window embeddings, and then aggregates the per-window edge predictions.
SAM-Road can predict the entire road network graph for arbitrarily large regions by operating in a sliding-window manner, as shown in Figure  3. The predictions within each window can be aggregated to improve accuracy. This is a common practice in computer vision applications for remote sensing, as fusing multiple observations can effectively suppress noise. For SAM-Road, this applies to both geometry and topology.
For the geometry, the per-window masks are fused to a large mask before vertex extraction, where each pixel value is the sum of all observed probabilities divided by the time it is observed. The NMS process is run on the fused global mask to get the global graph vertices.
@@
It’s also worth noting that the per-window inferences are completely independent of each other and can be done fully in parallel. This enables SAM-Road to be significantly faster (See Table  2) than the state-of-the-art RNGDet++ that reconstructs the graph in an auto-regressive manner. The ease of multi-window aggregation for quality refinement, akin to dense semantic segmentation, is also uncommon for typical graph-based methods. SAM-Road can flexibly trade-off between speed and accuracy, by varying the stride size in sliding-window inference, as shown in Table  3.
@@ 4.2 Metrics
We also utilize Average Path Length Similarity (APLS) [9] to evaluate the topological correctness. For a random vertex pair $(v_{1},v_{2})$ on the ground truth and their correspondences in the prediction $(\hat{v_{1}},\hat{v_{2}})$, we evaluate the model by comparing the shortest distance between $(v_{1},v_{2})$ and between $(\hat{v_{1}},\hat{v_{2}})$. Smaller distance difference indicates high topological similarity.
@@ 4.3 Implementation Details
At inference time, we use 16x16 sliding window inference for the main results. To determine the threshold for the binary classifiers (intersection, road, edge connection), we find the threshold that gives the highest F1 score on the validation set. Note that this is just for isolating away the effect of threshold choice in the experiments, and is not critical for SAM-Road performance, as evidenced by the result that just uses 0.5 for everything in Table  4.
@@ 4.4 Evaluating Road Network Prediction
Figure 4: The visualized road network graph predictions of SAM-Roadn and two baseline methods. Overall, SAM-Road generates highly accurate predictions. The circles highlight especially challenging spots: in the first area, SAM-Road correctly predicts the overpass structure. In the second one, SAM-Road gives superior results for the parallel freeways. The third spot shows an irregular intersection where the two baselines fail.
@@
Table 1: Comparison of the 8 prediction methods on different datasets. Our SAM-Road achieved the highest precision of $90.47\%$ on City-scale and $93.03\%$ on SpaceNet. It also shows the highest APLS metric of on both sets. Overall the graph accuracy is among the very top.
Qualitative results of SAM-Road predicting large-scale road network graphs can be found in Figure  4. The results are shown side-by-side with two baselines and the ground-truths. Overall, SAM-Road predicts highly accurate road networks even under very challenging circumstances, e.g. many blocks and intersections in dense urban areas, curvy roads with irregular shapes, overpasses, and multi-lane highways.
We benchmark SAM-Road on City-scale and SpaceNet benchmarks against other methods, quantitative results are shown in Table 1. We compare several baselines, including segmentation-based (Seg-UNet, Seg-DRM, Seg-Improved, Seg-DLA, Sat2Graph) and graph-based (RoadTracer, RNGDet, RNGDet++). The TOPO metric, which evaluates local graph structure similarity, is on par with state-of-the-art, RNGDet++, despite that SAM-Road has a much simpler structure without any bells and whistles. The APLS metric of SAM-Road achieves a new state-of-the-art. APLS evaluates the similarity of the shortest path that connects points A and B in two graphs, capturing long-range topological and geometrical structure - this indicates the effectiveness of our transformer-based topology decoder and graph representation.
Such performance should largely be attributed to SAM features, the highly powerful foundational vision model. As shown in 1, the predicted masks show high quality with little noise. This enables precise geometry prediction. The SAM image features are also informative vertex embeddings containing rich semantic meanings, as evident in the accurate topology predictions.
@@ 4.5 Speed and accuracy trade-off
SAM-Road is also highly efficient, thanks to its parallelized inference and it doesn’t require complex CPU-heavy post-processing heuristics. We measure the inference time to produce the complete graphs for the test sets of both datasets. The main results use $16\times 16$ windows and are already 40×40$\times$40 × faster than RNGDet++ on the City-scale dataset, and 10×10$\times$10 × faster on the SpaceNet dataset, as shown in 2. As mentioned in Section 3.6, SAM-Road can trade accuracy for more speed by specifying the sliding windows. Table 3 shows the result such trade-off. Using fewer windows can further provide 2×2$\times$2 × to 4×4$\times$4 × speed-up, with a minor accuracy drop.

### v2

@@ 1 Introduction
Figure 1: SAM-Road effectively predicts accurate road network graphs for dense urban regions, including roads with complex and irregular shapes, bridges, and multi-lane freeways. The corresponding segmentation masks are sharp and clear.
@@
Road network graphs are spatial representations of the structure and layout of road networks. They are typically stored in a vectorized format [20], consisting of vertices and edges. The vertices may represent intersections, and edges could stand for road segments. Large-scale road network graphs are vital for various applications: they enable navigation systems like Google Maps to determine optimal routes, assist in path planning for autonomous vehicles [18, 67], and help city planners in traffic analysis and optimization [4], to name a few. These applications call for accurate and efficient methods to automatically create such graphs, as they require scaling to huge regions and even near-continuous updating [10], which are astoundingly expensive when manually done. Therefore, systems for automatically generating such maps have tremendous application value and are under active research.
@@
Recently, the rapid growth of foundational models [2, 45, 52, 46] showcased their impressive capabilities. These models, which leverage flexible, high-capacity, and scalable architectures such as Transformers [53], are pre-trained through effective self-supervision [22] methods and unprecedentedly large datasets. This endows them with robust semantic reasoning and generalization. Segment Anything Model (SAM) [29] is such a foundational vision model. Built on the powerful ViT [16] architecture and trained with billions of images and masks, it demonstrates unparalleled semantic segmentation capabilities. This raises intriguing questions: How can SAM be applied to the prediction of road network graphs from satellite images, and how good can it be?
@@
In this work, we answer these questions by introducing the SAM-Road model, which adapts the SAM for generating large-scale, vectorized road network graphs. Incorporating domain knowledge from previous research in satellite mapping, we divide the problem into two main components: geometry prediction and topology reasoning.
@@
We model graph geometry with a set of 2D vertices that, when densely sampled, accurately reflect the graph’s overall shape. The SAM-Road model first predicts dense segmentation masks to indicate the likelihood of road elements such as lane segments and intersections, then it employs simple non-maximum suppression to convert the pixels into vertices of the desired density. Leveraging the inherent semantic segmentation capabilities of SAM, this method can effectively capture highly complex shapes (see Figure 1), which are common in dense urban areas.
@@
A notable challenge for segmentation-based mapping approaches is the difficulty of inferring topology from dense imagery. This branch of methods often relied on slow, complex and error-prone post-processing heuristics. Inspired by recent advances in graph learning [14, 49], we developed a transformer-based graph neural network as the second stage of our model. This network focuses on predicting the local subgraph around each vertex and determining connectivity with nearby vertices to establish the overall graph topology. It utilizes relative vertex positions and image embeddings from the SAM backbone to guide its predictions.
@@
Despite its straightforward design, SAM-Road achieves accuracy comparable to more complex state-of-the-art systems on two widely recognized satellite mapping datasets: City-scale [23] and SpaceNet [17]. Moreover, for large spatial areas spanning multi-square kilometers, its architecture supports high degrees of parallelism and rapid GPU inference, achieving speeds up to 80 times faster than existing methods. We hope that this work will inspire further exploration of foundational vision models in remote sensing and graph learning tasks.
@@ 2 Related Works
2.1 SAM and Its Applications
In 2023, the Segment Anything paper [29] proposed a foundational model for image segmentation, capable of generalizing across various tasks, enabling impressive zero-shot transfer capabilities. Through fine-tuning or direct adoption, SAM has been used in object detection [36], image inpainting [62], segmentation of medical images [39, 28, 55, 64], and remote sensing tasks [11]. Existing adaptations of SAM in remote sensing have focused more on simple segmentation and have not yet been applied to the production of road network graphs.
@@ 2.2 Road Network Graph Prediction
Segmentation-based methods [40, 6, 23] treat the task as a dense mask prediction. They represent the road network graph structure through one or more images, each detailing aspects such as road existence, intersections, orientation [6], and connectivity [23]. Post-processing heuristics, such as thinning [13, 66] and path-finding [30], are then employed to extract the vectorized graph structure. Benefits of this approach include 1) the ability of segmentation masks to represent complex geometries as a bottom-up volumetric representation [48], and 2) ease of parallel patch-wise inference for large areas, and subsequent result aggregation for refinement. However, the challenge of topology prediction persists: handcrafted heuristics often fail with poor mask quality; even with high-quality masks, deriving topology from them remains ill-formed. There exist no universal heuristics for all complex road structures, like multi-way intersections, multi-lane highways, and overpasses. Moreover, the heuristic tends to rely on CPU-intensive logic, which often becomes the inference speed bottleneck.
@@
Our method combines the advantages of segmentation-based and graph-based approaches. It harnesses the exceptional capabilities of SAM to generate a high-quality mask for geometry prediction, and uses a transformer-based graph neural network to directly produce graph structures without handcrafted post-processing heuristics.
@@ 2.3 Graph Representation and Learning
Graph representation and learning [21] involves mapping data to graph structures and applying learning algorithms to understand complex relationships within. Significant advancements have been made in this area with the development of Graph Neural Networks (GNNs) [56], Graph Convolutional Networks (GCNs) [31], and Transformers adapted for graph data [14]. Entities with rich structures can be represented as graphs and predicted by deep nets, such as scene graphs [19], human keypoints [34, 57], meshes [43], and in our case, road networks. The goal is to predict whether a graph edge (road segment) exists between a pair of nodes (vertices). For this type of task, GCN is a suitable architecture choice, as they offer powerful mechanisms for aggregating local subgraph information and understanding node relationships. With multiple layers, long-range dependencies can be captured too. In SAM-Road, we adopt Transformers as a special form of GCN: their self-attention mechanism has a simple form and can automatically select the most relevant context [3, 53, 18, 50] without any preset structure.
@@ 3.1 Overall Architecture
Figure 2: The architecture of our approach, SAM-Road. It contains an image encoder taken from the pre-trained SAM [29], a geometry decoder, and a topology decoder. It directly predicts vectorized graph vertices (yellow) and edges (orange) from an input RGB satellite imagery. Better zoom-in and view in color.
The overall structure of SAM-Road is shown in Figure  2. It contains an image encoder taken from the pre-trained SAM [29], a geometry decoder, and a topology decoder. The model takes as input an RGB satellite imagery. First, the image encoder produces the image feature embeddings. Then, the geometry decoder predicts the per-pixel existence probability, for both roads and intersections. The set of graph vertices $\mathbf{V}\{v_{i}\in\mathbb{R}^{2}\}$ representing 2D locations is extracted from these masks with a simple non-maximum suppression process, detailed in Algorithm 1. Given the predicted vertices, the topology decoder goes over each of them and determines whether it should connect to its nearby vertices within a given radius $R_{\text{nbr}}$ , given its local context. For an edge $(v_{i},v_{j})$, it predicts the probability that it exists. One edge may be predicted more than once, its final score will be the average. Eventually, the road network graph $\mathbf{G}$ is predicted as the sets of vertex $\mathbf{V}$ and edges $\mathbf{E}$
@@ 3.3 Geometry Decoder
We predict masks for both intersections and roads for more accurate graph structures at intersections. If only the road mask existed, there would be no guarantee that the center point of an intersection would be kept, producing error patterns like Figure  6. To mitigate this: 1) Vertices are extracted from both masks with the same NMS algorithm. 2) The two sets of vertices are joined, with all intersection vertices assigned a higher score than any road vertices. 3) The joined set is then NMS-ed again to produce the final result. This ensures intersection points are kept as much as possible.
@@ 3.4 Topology Decoder
The connection here is defined as ”whether two vertices are immediate neighbors on the graph”. That is, imagine a breadth-first-search on the road network graph from the source vertex, which stops expanding whenever a) it hits a target vertex or b) the depth (search radius) exceeds $R_{\text{nbr}}$ a target vertex is only connected to the source if it’s visited by the search. This is further illustrated in Figure  3.
@@ 3.5 Label Generation
Mask Labels. For road mask labels, we rasterize the ground-truth road lines, by drawing each edge as a line segment, with a width of 3 pixels. The pixels covered by the line segments are set to 1, and others are 0. For intersection labels, we find all the graph vertices with a degree not equal to 2 and render them as circles with a radius of 3 pixels. This is partially inspired by the OpenPose [8] work which represents human keypoint graphs as heatmaps.
Topology Labels. During training, we don’t run the vertex extraction process. The topology decoder is trained in a teacher-forcing [54] manner, where the vertices being asked are not from model prediction, but sampled from ground-truth road network graphs to emulate the predictions. This is done by first subdividing the ground-truth graph and then running the same NMS procedure as the inference stage. To emulate various NMS results, a uniform random score is assigned to each subdivision vertex.
@@ 3.6 Sliding-window Inference for Large Regions
Figure 4: SAM-Road can predict the entire road network graph for arbitrarily large regions by operating in a sliding-window manner. 0-3 represent 4 overlapping windows. It first extracts the global nodes, caches the per-window embeddings, and then aggregates the per-window edge predictions.
SAM-Road can predict the entire road network graph for arbitrarily large regions by operating in a sliding-window manner, as shown in Figure  4. The predictions within each window can be aggregated to improve accuracy. Fusing multiple observations is a common practice in vision applications [30, 44, 15, 33] to effectively suppress noise. For SAM-Road, this applies to both geometry and topology.
For geometry, the per-window masks are fused to a large mask before vertex extraction, where each pixel value is the sum of all observed probabilities divided by the time it is observed. The NMS process is run on the fused global mask to get the global graph vertices.
@@
It’s also worth noting that the per-window inferences are completely independent of each other and can be done fully in parallel. This enables SAM-Road to be significantly faster (See Table  2) than the state-of-the-art RNGDet++ [60] that reconstructs the graph in an auto-regressive manner. The ease of multi-window aggregation for quality refinement, akin to dense semantic segmentation, is also uncommon for typical graph-based methods. SAM-Road can flexibly trade-off between speed and accuracy, by varying the stride size in sliding-window inference, as shown in Table  3.
@@ 4.2 Metrics
We also utilize APLS (Average Path Length Similarity) [17] to evaluate the topological correctness. For a random vertex pair $(v_{1},v_{2})$ on the ground truth and their correspondences in the prediction $(\hat{v_{1}},\hat{v_{2}})$, we evaluate the model by comparing the shortest distance between $(v_{1},v_{2})$ and between $(\hat{v_{1}},\hat{v_{2}})$. Smaller distance difference indicates high topological similarity.
@@ 4.3 Implementation Details
At inference time, we use 16x16 sliding window inference for the main results. To determine the threshold for the binary classifiers (intersection, road, edge connection), we find the threshold that gives the highest F1 score on the validation set. Note that this is just for isolating away the effect of threshold choice in the experiments, and is not critical for SAM-Road performance, as evidenced by the result that just uses 0.5 for everything in Table  4 (A vs H).
@@ 4.4 Evaluating Road Network Prediction
Figure 5: The visualized road network graph predictions of SAM-Road and two baseline methods. Better zoom-in and view in color. Overall, SAM-Road generates highly accurate predictions. The circles highlight especially challenging spots: in the first area, SAM-Road correctly predicts the overpass structure. In the second one, SAM-Road gives superior results for the parallel freeways. The third spot shows an irregular intersection where the two baselines fail.
@@
Table 1: Comparison with existing methods on different datasets. SAM-Road achieved the highest TOPO precision of $90.47\%$ on City-scale and $93.03\%$ on SpaceNet. It also shows the highest APLS metric of on both sets. Overall the graph accuracy is among the very top.
Qualitative results of SAM-Road predicting large-scale road network graphs can be found in Figure  5. The results are shown side-by-side with two baselines and the ground-truths. Overall, SAM-Road predicts highly accurate road networks even under very challenging circumstances, e.g. many blocks and intersections in dense urban areas, curvy roads with irregular shapes, overpasses, and multi-lane highways.
We benchmark SAM-Road on City-scale and SpaceNet benchmarks against other methods, quantitative results are shown in Table 1. We compare several baselines, including segmentation-based (Seg-UNet, Seg-DRM, Seg-Improved, Seg-DLA, Sat2Graph) and graph-based (RoadTracer, RNGDet, RNGDet++). The TOPO metric, which evaluates local graph structure similarity, is on par with state-of-the-art, RNGDet++, despite that SAM-Road has a much simpler structure. The APLS metric of SAM-Road achieves a new state-of-the-art. APLS captures long-range topological and geometrical structure - this indicates the effectiveness of our transformer-based topology decoder and graph representation.
Such performance should largely be attributed to SAM features, the highly powerful foundational vision model. As shown in Figure 1, the predicted masks are sharp and clear, enabling precise geometry prediction. The SAM image features are also informative vertex embeddings containing rich semantic meanings, as evident in the accurate topology predictions.
@@ 4.5 Speed and Accuracy Trade-off
SAM-Road is also highly efficient, thanks to its parallelized inference and it doesn’t require complex CPU-heavy post-processing heuristics. We measure the inference time to produce the complete graphs for the test sets of both datasets. The main results use $16\times 16$ windows and are already 40×40$\times$40 × faster than RNGDet++ on the City-scale dataset, and 10×10$\times$10 × faster on the SpaceNet dataset, as shown in Table 2. As mentioned in Section 3.6, SAM-Road can trade accuracy for more speed by sparsifying the sliding windows. Table 3 shows the result such trade-off. Using fewer windows can further provide 2×2$\times$2 × to 4×4$\times$4 × speed-up, with a minor accuracy drop.